OpenStudy (fettiwap1738):

Part A: Explain why the x-coordinates of the points where the graphs of the equations y = 4–x and y = 2x + 3 intersect are the solutions of the equation 4–x = 2x + 3. (4 points) Part B: Make tables to find the solution to 4–x = 2x + 3. Take the integer values of x between −3 and 3. (4 points) Part C: How can you solve the equation 4–x = 2x + 3 graphically? (2 points) there is no way you can explain this to me but i can help you with history if you give me answers

OpenStudy (texaschic101):

not sure how to explain this....you have the y's equal to each other...so set their values equal to each other giving you 4 - x = 2x + 3....and once solved, you will have x...you can then sub that value back into either of the original equations to find y. 4 - x = 2x + 3 4 - 3 = 2x + x 1 = 3x 1/3 = x <== this is your solution